# The dark matter content of Milky Way dwarf spheroidal galaxies: Draco, Sextans and Ursa Minor

### v1

@@ 3.1 Single-population model
where $\nu$ represents the tracer number density distribution and $\Phi$ represents the total gravitational potential.
To obtain unique solutions to the Jeans equations, we further assume that the velocity ellipsoid is aligned with the cylindrical coordinates, so the cross-term can be neglected, i.e., $\overline{v_{R}v_{z}}=0$. Besides, the velocity anisotropy, defined as $\beta_{z}=1-\overline{v^{2}_{z}}/\overline{v^{2}_{R}}$, is assumed to be constant. Moreover, to calculate the first velocity moments, a rotation parameter $\kappa$ is defined as $\overline{v_{\phi}}=\kappa|\overline{v^{2}_{\phi}}-\overline{v^{2}_{R}}|^{1/2}$, which we also assume to be constant. Note our $\kappa$ follows the modified definition of Zhu et al. (2016a).
The gravitational potential ($\Phi$) in Equation 1 and 2 is contributed by both luminous and dark matter. For dark matter, we adopt the generalized NFW (Navarro et al., 1995, 1996b, 1997) (gNFW) model profile
@@
where $\rho_{s}$ is the scale density, $r_{s}$ is the scale radius, $\gamma$ and $\alpha$ are inner and outer density slopes, respectively. In our analysis we consider spherical dark matter halos. Moreover, because the number of member stars in outskirts is much less, our constraints on the outer density slope, $\alpha$, are very weak, and thus we fix $\alpha$ to 4 throughout the paper. We choose $4$ instead of $3$ because the outskirts of dSphs have undergone tidal effects so their profiles become steeper. However, we have verified that our constraints on the inner density profiles are not sensitive to the chosen value of the outer slope, and the constraints on the inner density profiles do not change as well if we make $\alpha$ a free parameter (also see Wang et al., 2022).
@@
For the luminous matter, i.e., member stars, the gravitational potential is obtained by deprojecting the density map based on the inclination angle. The density map is estimated by multiplying the observed surface brightness map by the stellar mass-to-light ratio ($M_{\ast}/L_{\ast}$). Here $M_{\ast}/L_{\ast}$ is determined from the closest matched PARSEC isochrone (Bressan et al., 2012) [^4] http://stev.oapd.inaf.it/cmd. To obtain the surface brightness map, we create a projected image for each dSph using membership catalogs constructed in Section 2.4, with each member star weighted by its luminosity. To account for faint stars with magnitudes below the Gaia flux limit, we scale the surface brightness map up to bring the integrated luminosity equal to the total luminosity of each dSph from Muñoz et al. (2018).
@@
In addition to the dark and luminous potential, there is the tracer surface number density term ($\nu$) in Equations 1 and 2. In principle, $\nu$ can be deprojected from the surface number density maps of member stars with spectroscopic observations, which are used as tracers. Note, however, for each dSph only a subset of stars in the full membership catalog constructed in Section 2.4 have spectroscopic observations due to fiber collisions. Especially in central regions of dSphs, where the surface number density of stars is very high, the completeness fraction of spectroscopically observed stars versus targets is much lower (e.g. Ding, et al., submitted). Thus instead of using the surface number density maps of spectroscopically observed member stars, $\nu$ is deprojected from the surface number density maps of all target stars in the full membership catalog in Section 2.4. We do not include any scaling on the amplitude of $\nu$, as it cancels out on both sides of the Jeans Equations.
@@
We show the surface number density profiles of all target member stars (black squares) and spectroscopically observed member stars (black triangles) in concentric elliptical radial bins in Figure 3. For each dSph in the upper panels, the projected axis ratio $q^{\prime}$ is obtained from the surface number density map and remains the same for different elliptical bins. The dashed curves in the upper panels are the Multi-Gaussian Expansions to these profiles, which will be introduced later. In the bottom panels, the completeness fractions for the spectroscopically observed member stars with respect to the targets in each dSphs are also shown. It can be seen that the completeness fractions are lower in the denser central regions due to unassigned fibers. Here the completeness fractions for Draco are different from those in Ding, et al., submitted, since we combine DESI MWS data with those from Walker et al. (2023) and remove stars on the horizontal branch.
@@
Although the spatial distributions of spectroscopically observed member stars and the full member star sample differ, we assume they do not differ in the velocity moment profiles, so the dynamical modeling outcome is not affected due to the incompleteness in fiber assignments. Equivalently, we are assuming that the spectroscopically observed member stars at different projected radii to the dSph center are a random subset in velocity space of the full sample. This is a reasonable assumption if the fiber assignment does not depend on any velocity information. However, as we have seen in Figure 3, the tiling and fiber completeness fraction differ in the central regions and the outskirts. In addition, metal-rich stars have more concentrated radial distributions to the dSph center, and the kinematics of metal-rich and metal-poor stars can differ. The fiber assignments may depend on member star kinematics. Unfortunately, if such a selection bias exists, it is hard to incorporate corrections, because we do not have LOSV measurements for the full target sample. We choose to adopt the assumption for the main results of our current paper, with detailed discussions provided in Section 5 to discuss such selection effects.
@@
where $L_{j}$, $\sigma_{j}$ and $q_{j}^{\prime}$ are the total number density, dispersion and projected axis ratio of each Gaussian component $j$. The corresponding 3-dimensional tracer number density distribution $\nu(R,z)$ (in cylindrical coordinates) can be simply deprojected as
@@
Note that the inclination angle $i$ is defined as the angle between the normal vector of the dSph and the line of sight, where $i=0$ corresponds to a face-on orientation.
Since the surface number density map is noisy, in practice we fit 1-dimensional MGE to surface number density profiles in Figure 3 to obtain $L_{j}$ and $\sigma_{j}$ for each Gaussian component $j$ in Equation 4. This is done in bins of concentric elliptical isophotes with the same axis ratio and orientation. Then these 1-dimensional MGE can be extended to 2-dimensional MGE by adding the projected axis ratio $q_{j}^{\prime}$ to each Gaussian component. $q_{j}^{\prime}$ is the axis ratio of the entire dwarf galaxies. Here we are assuming that different Gaussian components have the same axis ratio, which is a reasonable assumption based on the projected map of the three dSphs. We show the 1-dimensional MGE components for three dSphs in Figure 3 as colored dashed curves, and the black solid curve denotes the summation of these MGE components. The parameters of each MGE component and for the three dSphs are shown in Table 3. In all three dSphs, black solid curves provide good matches to black squares, which are surface number density profiles of all targeted member stars.
In the end, the final solution of the velocity moments is the summation of the analytical solution of each MGE component. In principle, the velocity anisotropy, $\beta_{z}$, and the rotation parameter, $\kappa$, can vary with different MGE components, but in this paper we assume that they are the same for all MGE components.
@@
Object $j$ $L_{j}$ $\sigma_{j}$ $q_{j}^{\prime}$
$[\mathrm{10^{-4}pc^{-2}}]$ $[\mathrm{pc}]$
Draco 1 40.27 50.99 0.65
2 136.26 122.48 0.65
3 21.27 233.47 0.65
Sextans 1 10.65 253.38 0.70
2 7.26 494.93 0.70
UMi 1 11.45 109.15 0.49
2 53.41 256.90 0.49
3 1.40 345.72 0.49
4 6.08 517.06 0.49
Table 3: Parameters of MGE for surface number density profiles of three dSphs. For each Gaussian component $j$, $L_{j}$ is the total number density, $\sigma_{j}$ is the dispersion along the major axis, and $q_{j}^{\prime}$ is the projected axis ratio.
The above analytical solutions are first obtained in the intrinsic coordinates of the dSph, and then transformed to the observed frame to be compared with observational data. The coordinate system in the observed frame $(x^{\prime},y^{\prime},z^{\prime})$ is a left-handed system. The $x^{\prime}$-axis lies along the projected major axis of the observed image of the galaxy, which is required by the MGE approach, and the $y^{\prime}$-axis lies along the projected minor axis. The $z^{\prime}$-axis is aligned with the line-of-sight direction, pointing away from us. On the other hand, the coordinate system in the intrinsic frame $(x,y,z)$ is a right-handed system. The $x$-axis coincides with the $x^{\prime}$-axis of the coordinates in the observed frame, and two coordinate systems are linked via
@@
where $\sigma_{v_{x^{\prime}},i}$, $\sigma_{v_{y^{\prime}},i}$ and $\sigma_{v_{z^{\prime}},i}$ are observational errors of velocity components $v_{x^{\prime},i}$, $v_{y^{\prime},i}$ and $v_{z^{\prime},i}$, respectively. Here the observational errors of $v_{x^{\prime},i}$ and $v_{y^{\prime},i}$ are obtained from their Gaia proper motion uncertainties. At the distances of the three classical dSphs, the proper motion uncertainties are very large, so the major constraints are contributed by the LOSVs. We have checked that including or not including Gaia proper motions for dSph member stars lead to almost the same results. Based on a set of parameters $\Theta$, the model predicts a mean velocity vector $\boldsymbol{\mu}_{i}=(\overline{u_{x^{\prime},i}},\,\overline{u_{y^{\prime},i}},\,\overline{u_{z^{\prime},i}})$ with a covariance matrix
@@
at the projected position of this star. By assuming that the velocity distribution predicted by the model is a trivariate Gaussian with the mean velocity $\boldsymbol{\mu}_{i}$ and the covariance $\boldsymbol{C}_{i}$, the likelihood for each star can be expressed as
@@
Then the total likelihood, defined as the product of likelihoods for all $N$ stars, is
@@
    Scale density of the dark matter halo, $\rho_{s}$: uniform distribution over $-15\leq\log_{10}[\rho_{s}/(\mathrm{M_{\odot}}\,\mathrm{pc^{-3}})]\leq 15$;
@@
    Velocity anisotropy, $\beta_{z}$: uniform distribution over $-\infty<-\ln(1-\beta_{z})<+\infty$;
@@
    Rotation parameter, $\kappa$: uniform distribution over $-\infty<\kappa<+\infty$.
In particular, to improve the efficiency of exploring the parameter space for $\rho_{s}$ and $r_{s}$, we fit redefined parameters $d_{1}=\log_{10}\rho_{s}^{2}r_{s}^{3}$ and $d_{2}=\log_{10}\rho_{s}$ instead. In addition, the velocity anisotropy $\beta_{z}$ is symmetrically transformed as $\lambda=-\ln(1-\beta_{z})$.
@@ 3.2.1 Spatial probability
To calculate the spatial probability, we need to construct the surface number density map for each population. If the maps of two populations are left completely free, there may be too many parameters which are difficult for our model to constrain. To reduce the number of free parameters, we first construct two template maps, and the surface number density maps of two populations are assumed to be linear combinations of these two template maps.
To construct the template maps, we first apply a simple hard cut in metallicity to divide spectroscopically observed member stars into two template populations. This hard cut is chosen to maximize [^5] We have also tried different choices of hard cuts, such as the cuts that leave some gap in between the two populations, so that they are more separated. Our best constrained model profiles remain almost the same. the half-number radii [^6] The radius within which it includes half of the stars in the sample. difference between the two template populations after the division. Next we compute the surface number density maps of the two template populations defined by this hard cut, with the maps fixed as template maps. We call the two templates as red and blue. Note, however, the template maps created in this way are subject to incompleteness due to fiber collisions, and the completeness fraction is lower in more central regions where the surface number density of target member stars is higher. To correct for the incompleteness, we first calculate the completeness fraction map according to the ratios of the surface number density maps between spectroscopically observed member stars and the full sample of target stars in the membership catalog (see Section 2.4). We multiply the red and blue template maps by the inverse of the completeness fraction map to correct for the incompleteness. The completeness fraction map is the same for both templates, as without spectroscopic metallicity measurements for target stars without fiber assignments, it is impossible for us to divide them into two populations.
@@
We show the template maps for the three dSphs in Figure 4. The hard cuts in metallicity for the red and blue templates are $-2.10$, $-1.87$, $-2.07$ for Draco, Sextans and UMi, respectively. The red population has higher number densities in the inner region and lower densities in the outer region than the blue population. Thus, the hard cut in metallicity naturally leads to a more concentrated, metal-rich template population and a more extended, metal-poor template population in each dSph.
@@ 3.2.4 Joint probability
    Scale density of the dark matter halo, $\rho_{s}$: uniform distribution over $-15\leq\log_{10}[\rho_{s}/(\mathrm{M_{\odot}}\,\mathrm{pc^{-3}})]\leq 15$;
@@
    Velocity anisotropies for the metal-rich population, $\beta_{z}^{\mathrm{mr}}$, and for the metal-poor population, $\beta_{z}^{\mathrm{mp}}$: uniform distribution over $-5\leq-\ln(1-\beta_{z}^{\mathrm{mr}})\leq+5$, $-5\leq-\ln(1-\beta_{z}^{\mathrm{mp}})\leq+5$;
@@
    Rotation parameters for the metal-rich population, $\kappa^{\mathrm{mr}}$, and for the metal-poor population, $\kappa^{\mathrm{mp}}$: uniform distribution over $-\infty<\kappa^{\mathrm{mr}}<+\infty$, $-\infty<\kappa^{\mathrm{mp}}<+\infty$;
@@
    Mean of metallicity distributions for the metal-rich population, $Z_{0}^{\mathrm{mr}}$, and for the metal-poor population, $Z_{0}^{\mathrm{mp}}$: uniform distribution over $-\infty<Z_{0}^{\mathrm{mr}}<+\infty$, $-\infty<Z_{0}^{\mathrm{mp}}<+\infty$;

### v2

@@ 3.1 Single-population model
where $\nu$ represents the tracer number density distribution and $\Phi$ represents the total gravitational potential. Throughout this paper, we use $(R,\phi,z)$ and $(x,y,z)$ to describe the cylindrical polar coordinate system and the Cartesian coordinate system in the intrinsic frame of the system, where $R^{2}=x^{2}+y^{2}$ and $z$ is the symmetry axis.
To obtain unique solutions to the Jeans equations, we further assume that the velocity ellipsoid is aligned with the cylindrical coordinates, so the cross-term can be neglected, i.e., $\overline{v_{R}v_{z}}=0$. Besides, the velocity anisotropy, defined as $\beta_{z}=1-\overline{v^{2}_{z}}/\overline{v^{2}_{R}}$, is assumed as constant. Moreover, to calculate the first velocity moments, a rotation parameter $\kappa$ is defined as $\overline{v_{\phi}}=\kappa|\overline{v^{2}_{\phi}}-\overline{v^{2}_{R}}|^{1/2}$, which we also assume as constant. Our $\kappa$ follows the modified definition of Zhu et al. (2016a).
The gravitational potential ($\Phi$) in Equations 1 and 2 is contributed by both luminous and dark matter. For dark matter, we adopt the generalized NFW (Navarro et al., 1995, 1996b, 1997) (gNFW) model profile
@@
where $\rho_{s}$ is the scale density, $r_{s}$ is the scale radius, $\gamma$ and $\alpha$ are inner and outer density slopes, respectively. In our analysis we consider spherical dark matter halos. Although the spherical dark matter halo assumption may bias the inferred dark matter content (e.g. Pfenniger et al., 1994; Sylos Labini et al., 2025), we have checked that best constrained inner densities remain consistent well within $1\sigma$ if we consider the axisymmetric dark matter halo leaving its axis ratio as a free parameter. The best-fit axis ratios of dark matter halos for the three dSphs are slightly larger than 1 ($\sim 1-1.5$). Moreover, because there are fewer member stars in outskirts, the outer density slope, $\alpha$, is weakly constrained. Thus we fix $\alpha$ to 4. We choose $4$ instead of $3$ because the outskirts of dSphs have undergone tidal effects so the slopes may be steeper (e.g., Errani & Navarro, 2021), but we have verified that our constraints on the inner density profiles are not sensitive to the chosen outer slope, and the constraints on the inner density profiles do not change if $\alpha$ is a free parameter (also see Wang et al., 2022).
@@
For the luminous matter, the gravitational potential is obtained by deprojecting the density map based on the inclination angle. The density map is estimated by multiplying the observed surface brightness map by the stellar mass-to-light ratio ($M_{\ast}/L_{\ast}$). Here $M_{\ast}/L_{\ast}$ is determined from the closest matched PARSEC isochrone (Bressan et al., 2012) [^5] http://stev.oapd.inaf.it/cmd. To obtain the surface brightness map, we create a projected image for each dSph using membership catalogs constructed in Section 2.4, with each member star weighted by its luminosity. To account for faint stars with magnitudes below the Gaia flux limit, we scale the surface brightness map up to bring the integrated luminosity equal to the total luminosity of each dSph from Muñoz et al. (2018).
@@
In addition to the dark and luminous potential, there is the tracer surface number density term ($\nu$). $\nu$ can be deprojected from the surface number density maps of member stars with spectroscopic observations, which are used as tracers. However, only a subset of stars in the full membership catalog constructed in Section 2.4 have spectroscopic observations due to fiber collisions. Especially in central regions of dSphs, where the surface number density of stars is very high, the completeness fraction of spectroscopically observed stars versus targets is much lower (e.g. Ding, et al., submitted). Thus instead of using the surface number density maps of spectroscopically observed member stars, $\nu$ is deprojected from the surface number density maps of all target stars in the full membership catalog in Section 2.4. We do not include any scaling on the amplitude of $\nu$, as it cancels out on both sides of the Equations.
@@
We show the surface number density profiles of all target member stars (black squares) and spectroscopically observed member stars (black triangles) in concentric elliptical radial bins in Figure 3. The projected axis ratio $q^{\prime}$ is obtained from the surface number density map and remains the same for different elliptical bins. The different dashed curves in the upper panels are the Multi-Gaussian Expansions, which will be introduced later. In the bottom panels, the completeness fractions for the spectroscopically observed member stars with respect to the targets in each dSphs are also shown. It can be seen that the completeness fractions are lower in the denser central regions due to unassigned fibers. Here the completeness fractions for Draco are different from those in Ding, et al., submitted, since we combine DESI MWS data with those from Walker et al. (2023) and remove stars on the horizontal branch.
@@
Although the spatial distributions of spectroscopically observed member stars and the full member stars differ, we assume they do not differ in velocity moments, so the dynamical modeling outcome is not affected due to the incompleteness in fiber assignments. Equivalently, we are assuming that the spectroscopically observed member stars at different projected radii to the dSph center are a random subset in velocity space of the full sample. This is a reasonable assumption if the fiber assignment does not depend on any velocity information. However, as we have seen in Figure 3, the tiling and fiber completeness fractions differ in the central regions and the outskirts. In addition, metal-rich stars have more concentrated radial distributions to the dSph center, and the kinematics of metal-rich and metal-poor stars can differ. The fiber assignments may depend on member star kinematics. Unfortunately, if such a selection bias exists, it is hard to incorporate corrections, because we do not have LOSV measurements for the full target sample. We choose to adopt the assumption for the main results, with detailed discussions provided in Section 5 to discuss such selection effects.
@@
where $L_{j}$, $\sigma_{j}$ and $q_{j}^{\prime}$ are the total number density, dispersion and projected axis ratio of each Gaussian component $j$. Here $x^{\prime}$ and $y^{\prime}$ are the coordinates of the projected major and minor axes of the observed image of the galaxy. The corresponding 3-dimensional tracer number density distribution $\nu(R,z)$ (in cylindrical coordinates) can be simply deprojected as
@@
The inclination angle $i$ is defined as the angle between the symmetry axis of the dSph and the line of sight, where $i=0$ corresponds to a face-on orientation.
Since the surface number density map is noisy, in practice we fit 1-dimensional MGE to surface number density profiles in Figure 3 to obtain $L_{j}$ and $\sigma_{j}$ for each Gaussian component $j$ in Equation 4. This is done in bins of concentric elliptical isophotes with the same axis ratio and orientation. Then these 1-dimensional MGE is extended to 2-dimensional MGE by adding the projected axis ratio $q_{j}^{\prime}$ to each component. $q_{j}^{\prime}$ is the axis ratio of the entire dwarf galaxy. Here we assume that different Gaussian components have the same axis ratio, which is reasonable according to the actual projected maps. We show the 1-dimensional MGE components for three dSphs in Figure 3 as colored dashed curves, and the black solid curve denotes the summation of these MGE components. In all three dSphs, black solid curves provide good matches to black squares, which are surface number density profiles of targeted member stars.
The solution of the velocity moments is the summation of the analytical solution of each MGE component. In principle, the velocity anisotropy, $\beta_{z}$, and the rotation parameter, $\kappa$, can vary with different MGEs, but here we assume they are the same for all MGE components.
@@
The solutions are first obtained in the intrinsic frames of the dSphs, and transformed to the observed frame. The coordinate system in the observed frame $(x^{\prime},y^{\prime},z^{\prime})$ is a left-handed system. The $x^{\prime}$-axis lies along the projected major axis of the observed image of the galaxy, which is required by the MGE approach, and the $y^{\prime}$-axis lies along the projected minor axis. The $z^{\prime}$-axis is aligned with the line-of-sight direction, pointing away from us. On the other hand, the coordinate system in the intrinsic frame $(x,y,z)$ is a right-handed system. The $x$-axis coincides with the $x^{\prime}$-axis of the coordinates in the observed frame, and two coordinate systems are linked via
@@
where $\sigma_{v_{x^{\prime}},i}$, $\sigma_{v_{y^{\prime}},i}$ and $\sigma_{v_{z^{\prime}},i}$ are observational errors of velocity components $v_{x^{\prime},i}$, $v_{y^{\prime},i}$ and $v_{z^{\prime},i}$, respectively. Here the observational errors of $v_{x^{\prime},i}$ and $v_{y^{\prime},i}$ are obtained from their Gaia proper motion uncertainties. At the distances of the three classical dSphs, the proper motion uncertainties are very large, so the major constraints are contributed by the LOSVs. Including or not including Gaia proper motions for dSph member stars lead to almost the same results. Based on a set of parameters $\Theta$, the model predicts a mean velocity vector $\boldsymbol{\mu}_{i}=(\overline{u_{x^{\prime},i}},\,\overline{u_{y^{\prime},i}},\,\overline{u_{z^{\prime},i}})$ with a covariance matrix
@@
at the projected position of this star. By assuming that the velocity distribution predicted by the model is a trivariate Gaussian with the mean velocity $\boldsymbol{\mu}_{i}$ and the covariance $\boldsymbol{C}_{i}$, the likelihood for each star is
@@
The total likelihood is the product of $L_{i}$ for all $N$ stars
@@
    Scale density of the dark matter halo, $\rho_{s}$: uniform distribution over $15\leq\log_{10}[\rho_{s}/(\mathrm{M_{\odot}}\,\mathrm{pc^{-3}})]\leq 15$;
@@
    Velocity anisotropy, $\beta_{z}$: uniform distribution over $\infty<-\ln(1-\beta_{z})<+\infty$;
@@
    Rotation parameter, $\kappa$: uniform distribution over $\infty<\kappa<+\infty$.
To improve the efficiency of exploring the parameter space for $\rho_{s}$ and $r_{s}$, we fit redefined parameters $d_{1}=\log_{10}\rho_{s}^{2}r_{s}^{3}$ and $d_{2}=\log_{10}\rho_{s}$. Besides, the velocity anisotropy $\beta_{z}$ is symmetrically transformed as $\lambda=-\ln(1-\beta_{z})$.
@@ 3.2.1 Spatial probability
For the spatial probability, if the maps of two populations are completely free, there are too many parameters which are difficult to constrain. To reduce the number of free parameters, we first construct two template maps, and the surface number density maps of two populations are assumed to be linear combinations of these two template maps.
To construct the template maps, we first apply a simple hard cut in metallicity to divide spectroscopically observed member stars into two template populations. This hard cut is chosen to maximize the half-number radii difference between the two template populations after the division. Next we compute the surface number density maps of the two template populations defined by this hard cut. We call the two templates as red and blue. Note, however, the template maps created in this way are subject to incompleteness due to fiber collisions, and the completeness fraction is lower in more central regions where the surface number density of target member stars is higher. To correct for the incompleteness, we first calculate the completeness fraction map according to the ratios of the surface number density maps between spectroscopically observed member stars and the full sample of target stars in the membership catalog (see Section 2.4). We multiply the red and blue template maps by the inverse of the completeness fraction map to correct for the incompleteness. The completeness fraction map is the same for both templates, as it is impossible for us to divide targets into two populations.
@@
The template maps are shown in Figure 4. The hard cuts in metallicity are 2.10-2.10, 1.87-1.87, 2.07-2.07 for Draco, Sextans and UMi, respectively. The red population has higher number densities in the inner region and lower densities in the outer region than the blue population. Thus, the hard cut in metallicity naturally leads to a more concentrated, metal-rich template population and a more extended, metal-poor template population.
@@ 3.2.4 Joint probability
    Scale density of the dark matter halo, $\rho_{s}$: uniform distribution over $15\leq\log_{10}[\rho_{s}/(\mathrm{M_{\odot}}\,\mathrm{pc^{-3}})]\leq 15$;
@@
    Velocity anisotropies for the metal-rich population, $\beta_{z}^{\mathrm{mr}}$, and for the metal-poor population, $\beta_{z}^{\mathrm{mp}}$: uniform distribution over $5\leq-\ln(1-\beta_{z}^{\mathrm{mr}})\leq+5$, $5\leq-\ln(1-\beta_{z}^{\mathrm{mp}})\leq+5$;
@@
    Rotation parameters for the metal-rich population, $\kappa^{\mathrm{mr}}$, and for the metal-poor population, $\kappa^{\mathrm{mp}}$: uniform distribution over $\infty<\kappa^{\mathrm{mr}}<+\infty$, $\infty<\kappa^{\mathrm{mp}}<+\infty$;
@@
    Mean of metallicity distributions for the metal-rich population, $Z_{0}^{\mathrm{mr}}$, and for the metal-poor population, $Z_{0}^{\mathrm{mp}}$: uniform distribution over $\infty<Z_{0}^{\mathrm{mr}}<+\infty$, $\infty<Z_{0}^{\mathrm{mp}}<+\infty$;